OpenStudy (anonymous):

arccos[sin(5pi/4)] Evaluate w/o using a calculator. No decimal answers and use pi radians is necessary.

OpenStudy (turingtest):

what is sin 5pi/4 ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

It is \[ \sin( \pi +\frac \pi 4) \]

OpenStudy (anonymous):

i still don't understand how to solve it

OpenStudy (anonymous):

find \(\frac{5\pi}{4}\)on the unit circle on the last page of the attached cheat sheet the first coordinate is cosine, the second coordinate is sine once you find \(\sin(\frac{5\pi}{4})\) find the angle between 0 and \(\pi\) whose cosine has the same value

OpenStudy (turingtest):

haha I have the same link with the same label^

OpenStudy (anonymous):

would the answer be 3pi/4

OpenStudy (anonymous):

you should find the point to have coordinates \((-\frac{\sqrt{2}}{2},-\frac{\sqrt{2}}{2})\) so \(\cos(\frac{5\pi}{4})=-\frac{\sqrt{2}}{2}\) and \(\sin(\frac{5\pi}{4})=-\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

yes, it would be \(\frac{3\pi}{4}\)

OpenStudy (anonymous):

You're a really good helper ! thank you.

OpenStudy (anonymous):

@TuringTest paul's notes! i was going to make my own, until i found this one and thought 'why bother?'

OpenStudy (anonymous):

@jeessicaaahh it is a great cheat sheet ( i can say that because i did not make it) so use it to your benefit. if you lose it google "paul's notes" and you will find lots of useful stuff for trig and also for future classes